OpenStudy (iwanttogotostanford):

HELP ASAP ONE PRE_ALGEBRA QUESTION PLEASE:

OpenStudy (iwanttogotostanford):

@jim_thompson5910 ?

OpenStudy (prettygirl_shynice):

(x^2 + 8x + 4) / (x^2 - x - 6) Factor the bottom. (x^2 + 8x + 4) / ( (x - 3) (x + 2) ) Set the bottom equal to zero. x - 3 = 0 x = 3 x + 2 = 0 x = -2 Confirm the x values do not cause the top to be zero (it would then be a hole, still a discontinuity, but not a asymptote). (-2)^2 + 8 (-2) + 4 = -8 (3)^2 + 8(3) + 4 = 37 Neither do so x = -2, 3 are both vertical asymptotes, aka discontinuities. Will ·

OpenStudy (prettygirl_shynice):

u got it

OpenStudy (iwanttogotostanford):

so i did do it right?

OpenStudy (iwanttogotostanford):

@Prettygirl_shynice

OpenStudy (prettygirl_shynice):

yea im here

OpenStudy (iwanttogotostanford):

Did i get the right answer then?

OpenStudy (prettygirl_shynice):

do u think u did

OpenStudy (iwanttogotostanford):

yes

OpenStudy (prettygirl_shynice):

well yea then

OpenStudy (iwanttogotostanford):

i came for a confirmation though

OpenStudy (prettygirl_shynice):

i understnad what you saying but yea u did

OpenStudy (prettygirl_shynice):

good job

OpenStudy (iwanttogotostanford):

thank you so much!:_)

OpenStudy (prettygirl_shynice):

@iwanttogotostanford ur welcome

OpenStudy (prettygirl_shynice):

u need help with anymore